# Uni-ISP: Unifying the Learning of ISPs from Multiple Cameras

### v1

@@ 3.3.3 Overall Loss
where $\mathcal{L}_{\text{NRR}}$ is the additional neutral rendering regularization and $\lambda$ is the balance weight for it. The neutral rendering regularization guides the model to learn a virtual camera that performs standard color conversion between XYZ and sRGB color space when the device-aware embedding is given as a zero vector $\emptyset$:
@@
where $s(\cdot)$ and $s^{-1}(\cdot)$ is the sRGB-XYZ and XYZ-sRGB color conversion respectively. The neutral rendering regularization adds an anchor for users if they want to enhance or reduce the photographic appearance of a certain camera without inter/extrapolating with the device-ware embedding of another camera.
3.4 Novel Dataset
Although existing sRGB-RAW datasets allow for training models on both inverse and forward ISP tasks, there remains a need for datasets that contain sRGB-RAW pairs captured synchronously by multiple devices. Such datasets are essential for training models to handle cross-camera ISP tasks effectively, which are crucial for applications like photographic appearance transfer and inter/extrapolation.
To address this challenge, we collected a novel dataset named FiveCam, featuring synchronously captured sRGB-RAW pairs from five distinct camera models. This dataset encompasses 2,464 high-resolution (4K) raw and JPEG images representing approximately 500 diverse scenes. Cameras used in the FiveCam dataset include the Apple iPhone 14 Pro Max, Google Pixel 6 Pro, Huawei P40, Samsung Galaxy S20, and Xiaomi Mi 12. All cameras are synchronized using a programmed Bluetooth shutter to ensure consistent timing across all devices.
A preview of the FiveCam dataset, illustrated in Figure 4, showcasing its three scenes alongside the capture devices used. Scenes of our FiveCam dataset are richly varied, capturing both natural landscapes and urban environments under multiple lighting conditions, ranging from broad daylight to nighttime settings, and including both outdoor and indoor illuminance.
Additionally, we have created an sRGB-XYZ version of this dataset, where raw images from all cameras are processed using as-shot white balance, a linear demosaicing algorithm, and converted to the standard camera-to-XYZ color space. The XYZ images in this version retain their linearity, making them particularly beneficial for downstream tasks that require maintaining the raw image’s linearity.
@@ 3.4 Novel Dataset
4 Applications
Table 1: Quantitative results of inverse and forward ISP tasks in the multi-camera mixed test set. All models are trained and tested with data from five different cameras. Please refer to Section 3.4 for detailed model names for each camera. To ensure fairness, the CycleISP [40] is not included in the forward ISP test since it requires ground truth RGB input during the forward ISP process.
@@ 4 Applications
4.1 Inverse and Forward ISP
Settings.
We compare Uni-ISP with previous ISP learning methods, including CycleISP [40], CIE-XYZNet [1], InvISP [38], and ParamISP [18]. All methods are retrained on the FiveCam dataset and we train two versions of baseline methods. The first version is trained on the mixed dataset, where sRGB-XYZ image pairs of all five cameras are used for training. The second version is trained on every single camera separately. Our Uni-ISP only trains a single version on the mixed dataset, with solely the self-camera training objective.
Results.
For the first version, we show quantitative results in Table 1, and qualitative results in Figure 5 and Figure 6. By observing the images and difference maps, our Uni-ISP has less error compared to other methods in both two tasks, which is consistent with the numeric results shown in Table 1. For the second version, results in Figure 7 demonstrate our Uni-ISP still outperforms all previous methods when previous methods are trained on a single device separately, i.e., five models for each previous method. In contrast, for our Uni-ISP, we only need to train a single model.
4.2 Photographic Appearance Transfer
Settings.
Photographic appearance transfer from camera model $a$ to camera model $b$ can be defined as,
$I_{b}=h(g(I_{a},\mathcal{E}_{a}),\mathcal{E}_{b}),$ (9)
where $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$ indicate two distinct device-aware embeddings of two camera models. This process uses inverse module $g$ to inverse the image to XYZ space and then uses forward module $h$ to generate the final image $I_{b}$. Given the absence of prior work, we select a series of global and local color transform methods as baselines, including learnable global color transforms, learnable 3D LUTs (Look-Up Tables), learnable bilateral grids, HDRnet [13] and NILUT [8]. We input the camera label as the style code within HDRnet and NILUT for a fair comparison. For the first three methods, we train 25 sub-models to accommodate the various transfer mappings required by the 25 possible pairs from the five devices in FiveCam. Still, we only need to train a single Uni-ISP. We adopt PSNR and DISTS [11] for evaluation.
Results.
As shown in Table 2, our method achieves the best results in terms of the two metrics, demonstrating the effectiveness of our method. A visual comparison presented in Figure 8 further illustrates that our Uni-ISP achieves results much closer to the ground truth compared to previous methods like HDRnet and NILUT, indicating the superior performance of our model. Moreover, our model has good temporal stability without a specific design. Figure 9 displays frames from the original iPhone video alongside those transferred to emulate the photographic styles of Google and Xiaomi. Please refer to the supplementary materials for the video.
Table 2: Quantitative results of photographic appearance transfer. The best value of each column is highlighted in bold.
@@ Results.
Table 3: Source camera identification accuracy of CMIResNet [7], MSFFN [14], and our Uni-ISP.
@@
4.3 Photographic Appearance Inter/Extrapolation
Settings.
In addition to transferring the photographic appearance, Uni-ISP presents the ability to inter/extrapolate the photographic appearance between two cameras smoothly. Given the photo $I_{a}$ captured by camera $a$, $I_{a}$ is firstly inverse to XYZ image by $g$. Then, the XYZ image is encoded to be the bottleneck feature and separately interacts with device-aware embeddings $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$, producing intermediate features $F_{a}$ and $F_{b}$ right after the DEIM. Then, the manipulated photo $I_{a+b}$ can be obtained by:
$I_{a+b}=h_{decoder}((1-\alpha){F}_{a}+\alpha{F}_{b}),$ (10)
Results.
The results of interpolating and extrapolating photographic appearance are demonstrated in Figure 10. In the first scenario, the Samsung Galaxy S20 prefers a deeper shadow and the extrapolation leads to a more “Samsung” style image. The second scenario is a night scene. Since Apple smartphones are famous for their night photography mode, our model also produces brighter results when we ask the model to be more “Apple”, where the lake and the trees become clearer.
4.4 Zero-shot Image Forensics
We perform two forensics tasks: source camera identification and image slice detection, without training on these tasks.
4.4.1 Source Camera Identification
@@
During the inference of source camera identification, Uni-ISP performs the inverse and forward ISP tasks multiple times with the device-aware embeddings of candidate camera models. Then we measure the edit distance $D$ of each inference, and the predicted camera model is the one that leads to the minimal edit distance. We empirically choose the structural similarity (SSIM) [36] as the metric for evaluating editing distance $D$. We compare our model with the camera model identification residual neural network (CMIResNet) [7] and the multi-scale feature fusion network (MSFFN) [14]. These methods are trained on the same training set as the Uni-ISP model, whose training objectives are classification tasks on the five cameras. For testing, we collect an extra test set containing 60 sRGB photos taken with the same five cameras in our proposed dataset, but in different scenes.
Results.
Table 3 shows the numeric results of the source camera identification experiments on our dataset. Based on these results, our methods show higher accuracy over the classification-based methods for source camera identification, revealing a new possible perspective on this task.
4.4.2 Image Splice Detection
@@
Our Uni-ISP offers additional capability in image splice detection. We can slightly modify its inference process for source camera identification to support image splice detection without extra training. Suppose an image $I_{a}$ is known to be taken with camera $a$, but it is altered. We can predict a map to indicate the areas likely to be altered using the self-consistency of inverse and forward ISP learned by Uni-ISP. This can be viewed as a pixel-level task in image forensics. We demonstrate this capability to show the versatility of our Uni-ISP.
Results.
Figure 11 demonstrates the example of Uni-ISP’s ability in image slice detection, where we put some fake objects on photos taken with iPhone 14 Pro Max. We harmonize the pasted fake objects using Poisson image editing [28]. The detection results are the visualization of the SSIM map, where white areas refer to minor differences, i.e., small editing distance, and darker areas indicate the location that may not belong to the original photo. The results prove the versatile ability when we unify the learning process of multiple camera ISPs.
Figure 11: Visualization of image slice detection by Uni-ISP. A fake object is integrated into a real photo using Poisson image editing [28], with darker regions in the result indicating suspected modifications.
4.5 Analysis
@@
Learning the inverse and forward ISP benefits HDR rendering. To validate the HDR rendering quality with currently learned inverse and forward ISP models, we synthesize multi-exposure raw image stacks with digital gains of $[1,2.6,4.2]$ on our test set, render them into sRGB images with a fixed ISP pipeline, and use the exposure fusion to produce the HDR tone-mapped images as ground truths. Given an LDR sRGB image, we adopt models for inverse ISP to convert the LDR sRGB image into an XYZ image. Then, we synthesize a stack of frames with the same digital gains as the dataset and follow the same manner of the dataset to render them into an HDR tone-mapped image.
Figure 12 shows the visualized results and each model’s performance in PSNR on the HDR rendering task. Consistent with the inverse ISP performance shown in Table 1, our method also achieves the best quality in the HDR rendering task.
@@
We conduct ablation on FBC loss by replacing it with an L1 loss that minimizes differences between model predictions and wrapped images for training. As shown in the image and spectrum visualization of Figure 13, the model trained with FBC loss produces clearer results than the one trained without FBC loss. Numeric results are presented in the supplementary material, which also demonstrates the FBC loss effectively avoids the frequency bias in the wrapped dataset.
Figure 13: Ablation results on the FBC loss. With FBC loss, we effectively avoid our model to learn the frequency bias in our wrapped dataset. The spectrum visualization of the noted patch also demonstrates that the model trained with FBC loss produces a sharper image with stronger high-frequent components.
@@
To elucidate why Uni-ISP effectively handles seamless inter/extrapolation of photographic appearances across multiple cameras, we visualize the internal feature distributions of our model. Specifically, we execute both self-camera tasks (inverse and forward ISP) and cross-camera tasks (photographic appearance transfer) using our test set. During these tests, we capture the forward ISP module $h$’s intermediate bottleneck features $B$ before they are processed by the DEIM and $F_{x}$ for each camera $x\in\{a,b,c,d,e\}$ after they are processed by the DEIM. Here, $x\in\{a,b,c,d,e\}$ denotes the variable for the five cameras in our new dataset. The location of $B$ and $F_{x}$ are noted in Figure 2. We utilize UMAP [26] to project these high-dimensional features into a 2D space, as shown in Figure 14. This visualization clearly delineates the transformation dynamics of features, with distinct clusters emerging for each camera model. This clustering visually confirms that the process of interpolating or extrapolating photographic appearances mimics sampling from a well-defined manifold of camera ISP behaviors. Thanks to the DEIM, which transforms features from an indistinct subspace (points noted by triangles) into a clearly delineable manifold (points noted by circles and squares), our model facilitates smooth inter/extrapolation between the photographic appearances of different camera models, as depicted in Figure 10.
@@
Given a few sRGB-Raw samples of an unseen camera, our Uni-ISP has the extension ability to perform well in inverse and forward ISP learning of that new camera by only training a new device-aware embedding while keeping the rest of the model frozen. We conduct the few-shot extension experiments on the S7 ISP dataset [32] to validate the few-shot extension ability of Uni-ISP. During training, we load the weights of the previously trained model in Section 4.1 and only optimize the newly created device-aware embedding for this unseen camera, Samsung Galaxy S7. For comparison, we select the ParamISP [18] since it outperforms previous methods in most cases. To simulate the scenario where only limited training data is available, we randomly select 10 scenes of the whole S7 dataset for training and the other 100 scenes for testing the generalizability. All models are trained for 50 epochs for convergence. Tests are conducted on the checkpoint of the 50th epoch. Table 4 shows the results of few-shot extension experiments with only 10 samples for learning this unseen camera model’s inverse and forward ISP.
As we can see, our Uni-ISP outperforms the ParamISP [18] in forward ISP when only 10 training samples are provided, which demonstrates the value of this extension ability in extreme scenarios of lacking adequate training datasets.
Table 4: Few-shot extension experiments on S7 ISP dataset [32] with only 10 samples for learning this unseen camera model’s inverse and forward ISP. We compare performances of the extension version of Uni-ISP (Uni-ISP-Ext) with the trained-from-scratch version of ParamISP [18] since the ParamISP outperforms other prior models in most cases. Numbers of trained parameters are given in the Param. column. The best performance for each few-shot sample scenario is noted in bold.
Method Param. Inverse ISP Forward ISP
PSNR ($\uparrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$)
ParamISP 700K 23.907 0.8305 21.439 0.8489
Uni-ISP-Ext 256 25.307 0.8597 22.798 0.8590
5 Conclusion
In this study, we presented the Uni-ISP, which can concurrently learn inverse and forward ISP behaviors of multiple cameras. The Uni-ISP harnesses the commonalities and distinctions across different cameras, enhancing the performance of ISP tasks beyond that of models trained on single-camera or mixed datasets. Additionally, Uni-ISP introduces novel applications for learned ISPs, including photographic appearance editing and zero-shot camera source forensics.
In future work, we plan to explore the interaction between the device-specific components and the backbone network by introducing more domain knowledge from ISP designs. Furthermore, considering the complexity of inverse ISP tasks, which can often be highly ill-posed, investigating the use of generative priors to maintain image fidelity presents an intriguing avenue for future research.

### v2

@@ III-C3 Overall Loss
where $\mathcal{L}_{\text{NRR}}$ is the additional neutral rendering regularization and $\lambda$ is the balance weight for it. The neutral rendering regularization guides the model to learn a virtual camera that performs standard color conversion between the XYZ and sRGB color spaces when the device-aware embedding is given as a zero vector $\emptyset$:
@@
where $s(\cdot)$ and $s^{-1}(\cdot)$ are the sRGB-XYZ and XYZ-sRGB color conversion, respectively. The neutral rendering regularization adds an anchor for users if they want to enhance or reduce the photographic appearance of a certain camera without inter/extrapolating with the device-ware embedding of another camera.
III-D Novel Dataset
Although existing sRGB-RAW datasets allow training models on both inverse and forward ISP tasks, there is still a need for datasets that contain synchronously captured sRGB-RAW pairs by multiple devices. Such datasets are essential for training models to handle cross-camera ISP tasks effectively, which are crucial for applications like photographic appearance transfer and inter/extrapolation.
To address this challenge, we collected a novel dataset named FiveCam, featuring synchronously captured sRGB-RAW pairs from five different camera models. This dataset encompasses 2,464 high-resolution (4K) raw and JPEG images representing approximately 500 diverse scenes. Cameras used in the FiveCam dataset include the Apple iPhone 14 Pro Max, Google Pixel 6 Pro, Huawei P40, Samsung Galaxy S20, and Xiaomi Mi 12. All cameras are synchronized using a programmed Bluetooth shutter to ensure consistent timing across all devices.
A preview of the FiveCam dataset, illustrated in Figure 5, showcasing its three scenes along with the capture devices used. The scenes in our FiveCam dataset are richly varied, capturing both natural landscapes and urban environments under multiple lighting conditions, ranging from broad daylight to nighttime settings, and including both outdoor and indoor illumination.
Additionally, we have created an sRGB-XYZ version of this dataset, where raw images from all cameras are processed using as-shot white balance, a linear demosaicing algorithm, and converted to the standard camera-to-XYZ color space. XYZ images in this version retain their linearity, making them particularly beneficial for downstream tasks that require maintaining the linearity of raw images.
@@ III-D Novel Dataset
IV Applications
TABLE I: Quantitative results of inverse and forward ISP tasks in the multi-camera mixed test set. All models are trained and tested with data from five different cameras. Please refer to Section III-D for detailed model names for each camera. To ensure fairness, the CycleISP [41] is not included in the forward ISP test since it requires ground truth RGB input during the forward ISP process.
@@ IV Applications
IV-A Inverse and Forward ISP
Settings
In the context of ISP learning, in addition to the forward ISP, inverse ISP enables raw-domain enhancements (e.g., HDR rendering, deblurring) from sRGB inputs when RAW data is unavailable, consistent with prior works [18, 39, 1]. We compare Uni-ISP with previous ISP learning methods in both inverse and forward ISP learning, including CycleISP [41], CIE-XYZNet [1], InvISP [39], and ParamISP [18]. All methods are retrained on the FiveCam dataset and we train two versions of baseline methods. The first version is trained on the mixed dataset, where sRGB-XYZ image pairs of all five cameras are used for training. The second version is trained on every single camera separately. Our Uni-ISP only trains a single version on the mixed dataset, with solely the self-camera training objective.
Results
For the first version, we show quantitative results in Table I, and qualitative results in Figure 6 and Figure 7. By observing the images and difference maps, our Uni-ISP has less error compared to other methods in both two tasks, which is consistent with the numeric results shown in Table I. For the second version, results in Figure 8 demonstrate our Uni-ISP still outperforms all previous methods even when previous methods are trained on a single device separately, i.e., five models for each previous method. In contrast, for our Uni-ISP, we only need to train a single model.
TABLE II: Quantitative results of photographic appearance transfer. The best value of each column is highlighted in bold.
@@ Results
IV-B Photographic Appearance Transfer
Settings
Photographic appearance transfer from the camera model $a$ to the camera model $b$ can be defined as:
$I_{b}=h(g(I_{a},\mathcal{E}_{a}),\mathcal{E}_{b}),$ (9)
where $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$ indicate two distinct device-aware embeddings of two camera models. This process uses the inverse module $g$ to inverse the image to the XYZ space and then uses forward module $h$ to generate the final image $I_{b}$. Given the absence of previous work, we select a series of global and local color transform methods as baselines, including learnable global color transforms, learnable 3D LUTs (Look-Up Tables), learnable bilateral grids, HDRnet [13] and NILUT [8]. We input the camera label as the style code within HDRnet and NILUT for a fair comparison. For the first three methods, we train 25 submodels to accommodate the various transfer mappings required by the 25 possible pairs of the five devices in FiveCam. Still, we only need to train one Uni-ISP. We adopt PSNR and DISTS [11] for evaluation.
Results
As shown in Table II, our method achieves the best results in terms of the two metrics, demonstrating the effectiveness of our method. We adopt DISTS to assess perceptual quality on unaligned original images, leveraging its robustness to structural differences. PSNR is computed on warped ground truth images ($I_{b}^{w}$), excluding occlusion areas identified by the optical flow method, to measure pixel-level accuracy in aligned regions. A visual comparison presented in Figure 9 further illustrates that our Uni-ISP achieves results much closer to the ground truth compared to previous methods such as HDRnet and NILUT, indicating the superior performance of our model. Moreover, our model has good temporal stability without a specific design. Figure 10 displays frames from the original iPhone video along with those transferred to emulate the photographic styles of Google and Xiaomi. Please refer to the supplementary materials for more details.
IV-C Photographic Appearance Inter/Extrapolation
Settings
In addition to transferring the photographic appearance, Uni-ISP presents the ability to smoothly inter/extrapolate the photographic appearance between two cameras. Given the photo $I_{a}$ captured by the camera $a$, $I_{a}$ is firstly inverse to the XYZ image by $g$. Then, the XYZ image is encoded as the bottleneck feature and interacts separately with device-aware embeddings $\mathcal{E}_{a}$ and $\mathcal{E}_{b}$, producing intermediate features $F_{a}$ and $F_{b}$ right after DEIM. Then, the manipulated photo $I_{a+b}$ can be obtained by:
$I_{a+b}=h_{decoder}((1-\alpha){F}_{a}+\alpha{F}_{b}),$ (10)
Results
The results of interpolating and extrapolating photographic appearance are demonstrated in Figure 11. In the first scenario, the Samsung Galaxy S20 prefers a deeper shadow, and the extrapolation leads to a darker image. The second scenario is a night scene. Since Apple smartphones are famous for their night photography, our model also produces brighter results when we ask the model to be more “Apple” where the lake and the trees become clearer.
Figure 12: Image splice detection samples processed with generative region-wise editing for realistic tampering. Our method predicts a confidence map, with darker areas indicating likely alterations from the original capture.
TABLE III: Source camera identification accuracy of CMIResNet [7], MSFFN [14], and our Uni-ISP.
@@
IV-D Zero-shot Image Forensics
We perform two forensics tasks: source camera identification and image splice detection, without training on these tasks.
IV-D1 Source Camera Identification
@@
During the inference of the source camera identification, Uni-ISP performs the inverse and forward ISP tasks multiple times with the device-aware embeddings of candidate camera models. Then we measure the edit distance $D$ of each inference, and the predicted camera model is the one that leads to the minimal edit distance. We empirically choose structural similarity (SSIM) [37] as the metric to evaluate the editing distance $D$. We compare our model with the camera model identification residual neural network (CMIResNet) [7] and the multi-scale feature fusion network (MSFFN) [14]. These methods are trained on the same training set as the Uni-ISP model, whose training objectives are classification tasks on the five cameras. For testing, we use 60 sRGB photos randomly chosen from a newly collected sRGB-only test set of unseen scenes captured by the same five cameras.
Results
Table III shows the numeric results of the source camera identification experiments. Based on these results, our methods show superior performance over the conventional classification-based methods for source camera identification on our dataset, revealing a new possible perspective on this task.
IV-D2 Image Splice Detection
@@
Suppose that an image $I_{a}$ is taken with a camera $a$, but it is altered. We can predict a map to indicate the areas that are likely to be altered using the self-consistency of Uni-ISP. This can be viewed as a pixel-level task in image forensics. We perform this task to show the versatility of our Uni-ISP.
Results
Figure 12 showcases Uni-ISP’s capability in image splice detection, where we introduce tampered objects into photos captured by an iPhone 14 Pro Max. These objects are seamlessly integrated using generative editing tools for realistic harmonization. The detection results are visualized as difference maps, with white regions indicating minimal differences (i.e., small editing distances) and darker regions highlighting areas likely altered from the original image. These results underscore Uni-ISP’s robust performance in unifying the learning of multiple camera ISPs for versatile forensic applications.
IV-E Analysis
@@
Learning the forward and inverse ISP benefits HDR rendering. To validate the HDR rendering quality with currently learned inverse and forward ISP models, we synthesize multi-exposure raw image stacks with digital gains of $[1,2.6,4.2]$ on our test set, render them into sRGB images with a fixed ISP pipeline, and use the exposure fusion to produce the HDR tone-mapped images as ground truths. Given an LDR sRGB image, we adopt models for inverse ISP to convert the LDR sRGB image into an XYZ image. Then, we synthesize a stack of frames with the same digital gains as the dataset and follow the same manner of the dataset to render them into an HDR tone-mapped image.
Figure 13 shows the visualized results and the performance of each model in PSNR on the HDR rendering task. Consistent with the inverse ISP performance shown in Table I, our method also achieves the best quality in the HDR rendering task.
@@
Evaluation on Raw-Domain Deblurring
The inverse ISP ability of Uni-ISP also makes it useful in raw-domain deblurring. We evaluate such ability of the Uni-ISP and other inverse ISP methods on the RealBlur dataset [31]. In this evaluation, we assume that we only have access to sRGB JPEG images, but we still want to conduct deblurring in the raw domain. We adopt the Stripformer [35] pretrained on raw images as the raw-domain deblurring network.
TABLE IV: Numeric results of raw-domain deblurring using Stripformer [35] on the raw images produces by inverse ISP methods, with sRGB images given as input.
Inverse ISP Method PSNR ($\uparrow$) SSIM ($\uparrow$)
CycleISP [41] 36.482 0.9382
CIE-XYZ Net [1] 35.074 0.9619
InvISP [39] 38.173 0.9767
ParamISP [18] 40.495 0.9784
Uni-ISP 40.847 0.9799
Table IV shows the results of performing the raw-domain deblurring using inverse ISP methods when only sRGB images are available. Our Uni-ISP outperforms other inverse ISP methods in this raw-domain deblurring task, aligning with the inverse ISP accuracy we showed in Section IV-A.
@@
We perform ablation on FBC loss by replacing it with an L1 loss that minimizes differences between model predictions and warped images for training. As shown in the image and spectrum visualization of Figure 14, the model trained with FBC loss produces clearer results than the one trained without FBC loss. Numeric results are presented in the supplementary material, which also demonstrates that the FBC loss effectively avoids frequency bias in the warped dataset.
Figure 14: Ablation results on the FBC loss. With FBC loss, we effectively avoid our model to learn the frequency bias in our warped dataset. The spectrum visualization of the noted patch also demonstrates that the model trained with FBC loss produces a sharper image with stronger high-frequent components.
TABLE V: Ablation study on components in the model design of Uni-ISP (GFMB, LFEB, and DEIM). The evaluation is conducted in self-camera ISP tasks. The best performance is noted in bold.
Components Inverse ISP Forward ISP
GFMB LFEB DEIM PSNR SSIM PSNR SSIM
26.181 0.8098 21.359 0.8658
28.782 0.8768 25.806 0.9090
29.010 0.8832 25.878 0.9094
30.940 0.9097 25.701 0.9102
32.320 0.9354 28.340 0.9245
32.699 0.9396 29.151 0.9306
Ablation Study on Components of Uni-ISP
We conduct ablation experiments on the LEFB, GFMB, and DEIM of Uni-ISP. When the LFEB is disabled, we use an ordinary convolutional layer from vanilla U-Net as replacement. When the GFMB is disabled, it simply becomes a residual skip connection. When the DEIM is disabled, the bottleneck feature will be passed to the decoder stage directly from the encoder stage. The results of the ablation experiments on two self-camera ISP tasks, which involve learning the inverse and forward ISP, are presented in Table V.
The LEFB extracts multi-scale features while the GFMB provides global manipulation based on the EXIF metadata. Therefore, both components contribute to improving the performance of Uni-ISP. It is worth noting that the first row in Table V shows a significant performance drop without the DEIM. This indicates that Uni-ISP becomes a non-device-aware model as in all previous works, leading to a drastic performance drop in tasks involving multiple camera devices.
@@
To elucidate why Uni-ISP effectively handles seamless inter/extrapolation of photographic appearances across multiple cameras, we visualize the internal feature distributions of our model. Specifically, we execute both self-camera tasks (inverse and forward ISP) and cross-camera tasks (photographic appearance transfer) using our test set. During these tests, we capture the forward ISP module $h$’s intermediate bottleneck features $B$ before they are processed by the DEIM and $F_{x}$ for each camera $x\in\{a,b,c,d,e\}$ after they are processed by the DEIM. Here, $x\in\{a,b,c,d,e\}$ denotes the variable for the five cameras in our new dataset. The locations of $B$ and $F_{x}$ are noted in Figure 2. We utilize UMAP [26] to project these high-dimensional features into a 2D space, as shown in Figure 15. This visualization clearly delineates the transformation dynamics of features, with distinct clusters emerging for each camera model. This clustering visually confirms that the process of interpolating or extrapolating photographic appearances mimics sampling from a well-defined manifold of camera ISP behaviors. Thanks to the DEIM, which transforms features from an indistinct subspace (points noted by triangles) into a clearly delineable manifold (points noted by circles and squares), our model facilitates smooth inter/extrapolation between the photographic appearances of different camera models, as depicted in Figure 11.
@@
TABLE VI: Few-shot extension experiments for learning unseen cameras’ inverse and forward ISP. Images are taken from the S7 ISP dataset [32] and our additional extension sub-set, covering mobile cameras including Samsung Galaxy S7, Apple iPhone 15 Pro, Apple iPhone 8 Plus, and SONY Xperia 5 II. The best performance for each camera is noted in bold.
Camera Method Inverse ISP Forward ISP
PSNR ($\uparrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$)
S7 ParamISP 23.907 0.8305 21.439 0.8489
Uni-ISP 25.307 0.8597 22.798 0.8590
15Pro ParamISP 24.281 0.5122 16.377 0.6266
Uni-ISP 25.301 0.7380 23.086 0.8429
8Plus ParamISP 14.952 0.3494 17.238 0.7046
Uni-ISP 24.922 0.8396 23.354 0.9251
Xperia 5 II ParamISP 13.802 0.1360 16.619 0.3985
Uni-ISP 19.848 0.6428 20.480 0.518
@@
The proposed Uni-ISP has the extension ability to perform well in the inverse and forward ISP learning of that new camera by only training a new device-aware embedding while keeping the rest of the model frozen. This requires a few sRGB-Raw pair samples captured by the target unseen camera. Adding a new camera requires collecting paired data for retraining, as modifying EXIF metadata alone is not sufficient to mimic the behavior of a new ISP.
We conduct the few-shot extension experiments on the S7 ISP dataset [32] and an additional small-scale dataset collected by ourselves using three additional cameras (iPhone 15 Pro, iPhone 8 Plus, and SONY Xperia 5 II) to validate the few-shot extension ability of Uni-ISP. During training, we load the weights of the previously trained model in Section IV-A and only optimize the newly created device-aware embedding for this unseen camera. For comparison, we select ParamISP [18] since it outperforms previous methods in most cases. Unlike our method, which only requires training new device-aware embeddings for unseen cameras, ParamISP requires full retraining to adapt to a new camera. Simply modifying the EXIF input of ParamISP is not enough to mimic an unseen camera’s ISP behavior. We randomly select 10 scenes taken by each camera for training and use the other 100 scenes for testing. All models are trained for 50 epochs for convergence. Tests are conducted at the checkpoint of the 50th epoch. Table VI shows the results of few-shot extension experiments to learn the inverse and forward ISP of unseen camera models. As we can see, our Uni-ISP outperforms ParamISP [18] when only few training samples are provided, which demonstrates the value of this extension ability in extreme scenarios with inadequate training datasets.
V Conclusion
In this study, we presented the Uni-ISP, which can concurrently learn inverse and forward ISP behaviors of multiple cameras. The Uni-ISP harnesses the commonalities and distinctions across different cameras, enhancing the performance of ISP tasks beyond that of models trained on single-camera or mixed datasets. Furthermore, Uni-ISP introduces novel applications for learned ISPs, including photographic appearance editing and zero-shot camera source forensics.
In future work, we plan to explore the interaction between device-specific components and the backbone network by introducing more domain knowledge from ISP designs. In addition, considering the complexity of inverse ISP tasks, which can often be highly ill-posed, investigating the use of generative priors to maintain image fidelity presents an intriguing avenue for future research.